Find the equation of integral surface to the differential equation
2y(z-3)p+(2x-z)q=y(2x-3)
1
Expert's answer
2020-10-05T17:43:35-0400

Given equation is 2y(z-3)p+(2x-z)q=y(2x-3), circle is x2+y2=2x,z=0

Langrange's auxiliary equations for equation are dx/(2y(z-3))=dy/(2x-z)=dz/(y(2x-3))

Taking the first and third fractions of it, (2x-3)dz-2(z-3)dz=0

Integrating, x2-3x-z2+6z=c1.c1 being an arbitrary constant.

Choosing 1/2,y,-1 as multipliers, each fraction of dx/(2y(z-3))=dy/(2x-z)=dz/(y(2x-3))=((1/2)dx+ydy-dz)/(y(z-3)+y(2x-3)-y(2x-3))=((1/2)dz+ydy-dz)/0

Hence (1/2)dx+ydy-dz=0 or dx+2ydy-2dz=0.

Integrating,x+y2-2z=c2,c2 being an arbitrary constant.

Now, the parametric equations of given circle are x=t, y=(2t-t2)1/2,z=0

substituting this values we have

t2-3t=c1 and 3t-t2=c2

Eliminating t from the above equations we have c1+c2=0

Substituting the values of c1 and c2 the desired integral surface is

x2-3x-z2+6z+x+y2-2z=0 or x2+y2-z2-2x+4z=0